OpenStudy (anonymous):

If a square has an area of 36 in2, how long is each side? Select the best answer from the choices provided. 1,296 in 12 in 24 in 6 in

OpenStudy (johnweldon1993):

Hint, square means 4 equal sides... And the area of a square = l times w So, since l and w are equal we have \[\large A = LW\] \[\large 36 = L^2\] \[\large L = \sqrt{36}\] \[\large L = ?\]

OpenStudy (anonymous):

6 ? @johnweldon1993

OpenStudy (johnweldon1993):

There ya go :)